# One Model to Forecast Them All

### v1

@@ I Introduction
Uncertainty is inherent in modern power and energy systems, intensified by the growing integration of renewable energy sources and demand response programs that require close monitoring of distribution grids. Reliable forecasting is critical to ensuring system reliability, efficient resource allocation, and operational affordability in these volatile environments. While point forecasting methods provide deterministic estimates, they fail to capture uncertainty. Probabilistic forecasting addresses this limitation by offering richer insights through probability distributions, confidence intervals, or quantiles, enabling better risk management and planning [1].
@@
Figure 1: The different utilization levels of probabilistic forecasting are illustrated in a two-dimensional space in representing average load in subsequent 12-hour intervals. First, the (possibly) complex forecasting distribution is approximated as a mixture of Gaussians. Then, a collection of samples is taken from it to form an ensemble. Finally, quantiles of each marginal are extracted using this ensemble.
@@
We adopt GUIDE-VAE [6], a conditional Variational Autoencoder designed for multi-entity and multivariate data, to address these challenges modelling. GUIDE-VAE combines deep generative modelling with two advanced features: Pattern Dictionary-based Covariance Composition (PDCC) and probabilistic entity embeddings. These features enable GUIDE-VAE to perform parallel probabilistic forecasting for individual entities using a single model, eliminating the inefficiency of training separate models. GUIDE-VAE also provides flexible forecasting outputs, from interpretable point estimates to full probability distributions, capturing uncertainty and temporal dependencies [7].
In this study, we focus on household electricity consumption as a challenging case study for multi-entity probabilistic forecasting. Household data is highly volatile and stochastic, making it an ideal testbed to demonstrate GUIDE-VAE’s scalability and modelling capabilities. The key contributions of this work are: (1) adapting GUIDE-VAE for day-ahead forecasting of mult-entity electricity consumption, (2) providing forecast outputs as distributions, samples, and quantiles, and (3) validating its effectiveness through experiments using household electricity consumption data.
@@ III Methodology
GUIDE-VAE [6] is a conditional VAE model that approximates the probability distribution model [^2] In order to optimize such a distribution model, all VAE-based models, including GUIDE-VAE, employ an amortized variational inference methodology. This requires an extra distribution model for the approximate posterior (an encoder network), which maps observable and conditioning variables into latent variables. Since this posterior distribution is essential only for the training of GUIDE-VAE and not significant for the proposed methodology, we refer interested readers to [3] and [6] for further details. $p_{\psi,\mathbf{U}}(\mathbf{x}|\mathbf{c})=\mathbb{E}_{p(\mathbf{z})}\left[p_{\psi,\mathbf{U}}(\mathbf{x}|\mathbf{z},\mathbf{c})\right]$. Here, $\mathbf{x}\in\mathbb{R}^{T}$, $\mathbf{c}\in\mathbb{R}^{C}$ and $\mathbf{z}\in\mathbb{R}^{Z}$ represent observable, conditioning and latent variables, respectively. This structure allows generating data points for a given condition $\mathbf{c}$ using ancestral sampling as $\hat{\mathbf{x}}\sim p_{\psi,\mathbf{U}}(\mathbf{x}|\hat{\mathbf{z}},\mathbf{c})$ where $\hat{\mathbf{z}}\sim p(\mathbf{z})$ as illustrated in Fig. 2. Here, the likelihood and prior distributions are respectively modelled as
@@
The parameterization function for the mean $\mu_{\psi}$ can be chosen without a constraint and often modelled using a neural network $f^{\mu}_{\psi}(\mathbf{z},\mathbf{c})$. However, the parametrization of the covariance requires satisfying positive definiteness, and common practice is to model it as a positive diagonal matrix. However, this choice results in an independence between the components of generated samples and hinders the modelling power.
@@ IV-A Dataset
This study uses a dataset of smart meter data collected from electricity consumers across 47 provinces in Spain, including homes, offices, and businesses [9]. The dataset spans from November 2014 to June 2022 and includes hourly electricity consumption records for 25,559 users. For the experiments, a subset was extracted focusing on data from Gipuzkoa, the province with the highest data density, between June 2021 and June 2022. Only users with at least one full year of continuous enrollment during this period were included, while those with consistently zero or negative consumption were excluded. The final dataset comprises $U$=6,830 users, each with $N$=365 daily profiles (24 hourly values per day, $T$=24), amounting to approximately 2.5 million records.
Since the data is zero-inflated, we used the same hyper-parameters to apply a zero-preserved log-normalization to each feature as described in [6]. After normalization, we reserved the last 73 days of all users’ records (20% of the whole dataset) as the testing set. At the same time, the remaining portion was split into training and validation sets with a ratio of 3:1. We used the validation set for early stopping and learning rate scheduling of every model we trained.
We conduct the experiments using two auxiliary conditioning features: months and weekdays. These conditions are derived from the data index $n$, transformed using a cyclic ($\sin$-$\cos$) encoding as proposed in [10], and associated with their respective data points $\mathbf{x}_{un}$. As for the user embeddings, we used the same hyper-parameter setting employed in [6].
@@ IV-B Baseline Model
$\mathbf{Y}=[\mathbf{y}_{i}]_{i=1}^{Q}=[[y^{(t)}_{i}]_{t=1}^{T}]_{i=1}^{Q}=f_{\phi}(\mathbf{x},\mathbf{c})$ (6)
where $y^{(t)}_{i}$ represents the $q^{(i)}$-th quantile value of the $t$-th feature for a given regressor $\mathbf{x}$ and contextual information $\mathbf{c}$ mapped using a neural network $f_{\phi}$. This model can be trained by minimizing the following loss function
@@
for a given dataset $\{(\mathbf{y}_{m},\mathbf{x}_{m},\mathbf{c}_{m})\}_{m=1}^{M}$ and predictions $\{[\mathbf{y}_{mi}]_{i=1}^{Q}=f_{\phi}(\mathbf{x}_{m},\mathbf{c}_{m})\}_{m=1}^{M}$. Note that this corresponds to the conventional quantile loss averaged over all output features and quantile levels. This setup can easily be converted to the day-ahead forecasting task at hand by replacing $\mathbf{y}_{m}$ with $\mathbf{x}_{un}$, $\mathbf{x}_{m}$ with $\mathbf{x}_{u(n-1)}$ and $\mathbf{c}_{m}$ with $\mathbf{c}_{un}$.
@@ IV-C Training
We applied the same training, neural network, and constraint settings on GUIDE-VAE as in [6], including features like the number of layers and neurons, learning rate, early stopping and learning rate schedule. Also, we set the size of the quantile regression baseline model equal to the GUIDE-VAE in terms of the number of learnable parameters to keep the comparison fair.
@@ IV-D1 Benchmarking Metrics
These metrics are used to test the performance of GUIDE-VAE against the quantile regression model. Since the quantile regression model does not yield an explicit probability distribution against which to compare, we adapted GUIDE-VAE to the quantile forecasting setting. For this, we took $S$ samples from $p_{\psi,\mathbf{U}}(\mathbf{x}_{un}|\mathbf{c}_{un})$ for each ($u$,$n$)-pair as $\hat{\mathbf{x}}_{uns}\sim p_{\psi,\mathbf{U}}(\mathbf{x}_{un}|\mathbf{z}_{s},\mathbf{c}_{un}),\forall s\in[S]$ where $\mathbf{z}_{s}\sim p(\mathbf{z})$. Recall that each condition $\mathbf{c}_{un}=\left[\mathbf{x}_{u(n-1)},\hat{\theta}_{u},\tilde{\mathbf{c}}_{n}\right]$ represents the look-back window and the contextual information required for forecasting $\mathbf{x}_{un}$. Then, we calculated the quantile values of each feature $x_{un}^{(t)}$ for given quantile levels $\mathbf{q}=[q^{(i)}]$ using these $S$ samples as
@@
Having access to quantile values for both forecasting models, we hereby enlist the benchmarking metrics and their definitions:
@@ IV-D2 Ablation Metrics
    $\text{ALL}=\frac{1}{UN_{\text{test}}}\sum_{u,n}^{U,N_{\text{test}}}\log p_{\psi,\mathbf{U}}(\mathbf{x}_{un}|\mathbf{x}_{u(n-1)},\tilde{\mathbf{c}}_{un})$ (11)
@@ VI Conclusion
This study introduced a probabilistic day-ahead forecasting framework based on GUIDE-VAE, a conditional VAE model designed for multi-entity datasets. GUIDE-VAE incorporates key features such as a pattern dictionary and entity-guided embeddings, enabling efficient modelling of complex data structures. Unlike traditional approaches, GUIDE-VAE supports scalable, entity-specific probabilistic forecasting with a single model, eliminating the inefficiency of training separate models for individual entities. Its flexible outputs range from interpretable point estimates to full probability distributions that capture uncertainty and temporal dependencies.
Household electricity consumption was used as a challenging case study due to its highly stochastic nature and large number of entities. Our results demonstrate that GUIDE-VAE significantly outperforms conventional quantile regression techniques across key metrics, including quantile loss, mean absolute deviation and interval-based scores. Additionally, the ablation study highlights the critical role of PDCC in capturing temporal correlations and improving forecast realism and accuracy across diverse settings.
In conclusion, GUIDE-VAE offers a powerful and generalizable framework for probabilistic forecasting tasks in energy systems and beyond. Its ability to handle multi-entity data, provide entity-specific forecasts, and deliver probabilistic outputs makes it a versatile tool for decision-makers aiming to improve risk management, resource allocation, and operational planning. Future work will focus on extending GUIDE-VAE to other multi-entity datasets, such as feeders and wind turbines, and exploring the use of its covariance structures for refining intra-day forecasts and uncovering deeper insights into temporal dependencies.

### v2

@@ I Introduction
Uncertainty is inherent in modern power and energy systems, intensified by the growing integration of renewable energy sources and demand response programs that require close monitoring of distribution grids. Reliable forecasting is critical for ensuring system reliability, efficient resource allocation, and operational affordability in these volatile environments. While point forecasting methods provide deterministic estimates, they fail to capture uncertainty. Probabilistic forecasting addresses this limitation by offering richer insights through probability distributions, confidence intervals, or quantiles, enabling better risk management and planning [1].
@@
Figure 1: The different utilization levels of probabilistic forecasting are illustrated in a two-dimensional space, representing average load in subsequent 12-hour intervals. First, the (possibly) complex forecasting distribution is approximated as a mixture of Gaussians. Then, a collection of samples is taken from it to form an ensemble. Finally, quantiles of each marginal are extracted using this ensemble.
@@
We adopt GUIDE-VAE [6], a conditional Variational Autoencoder designed for multi-entity and multivariate data, to address these challenges in modelling. GUIDE-VAE combines deep generative modelling with two advanced features: Pattern Dictionary-based Covariance Composition (PDCC) and probabilistic entity embeddings. These features enable GUIDE-VAE to perform parallel probabilistic forecasting for individual entities using a single model, eliminating the inefficiency of training separate models. GUIDE-VAE also provides flexible forecasting outputs, from interpretable point estimates to full probability distributions, capturing uncertainty and temporal dependencies [7].
In this study, we focus on household electricity consumption as a challenging case study for multi-entity probabilistic forecasting. Household data is highly volatile and stochastic, making it an ideal testbed to demonstrate GUIDE-VAE’s scalability and modelling capabilities. The key contributions of this work are: (1) adapting GUIDE-VAE for day-ahead forecasting of multi-entity electricity consumption, (2) providing forecast outputs as distributions, samples, and quantiles, and (3) validating its effectiveness through experiments using household electricity consumption data.
@@ III Methodology
GUIDE-VAE [6] is a conditional VAE model that approximates the probability distribution model [^2] In order to optimize such a distribution model, all VAE-based models, including GUIDE-VAE, employ an amortized variational inference methodology. This requires an extra distribution model for the approximate posterior (a decoder network), which maps observable and conditioning variables into latent variables. Since this posterior distribution is essential only for the training of GUIDE-VAE and not significant for the proposed methodology, we refer interested readers to [3] and [6] for further details. $p_{\psi,\mathbf{U}}(\mathbf{x}|\mathbf{c})=\mathbb{E}_{p(\mathbf{z})}\left[p_{\psi,\mathbf{U}}(\mathbf{x}|\mathbf{z},\mathbf{c})\right]$. Here, $\mathbf{x}\in\mathbb{R}^{T}$, $\mathbf{c}\in\mathbb{R}^{C}$ and $\mathbf{z}\in\mathbb{R}^{Z}$ represent observable, conditioning and latent variables, respectively. This structure allows generating data points for a given condition $\mathbf{c}$ using ancestral sampling as $\hat{\mathbf{x}}\sim p_{\psi,\mathbf{U}}(\mathbf{x}|\hat{\mathbf{z}},\mathbf{c})$ where $\hat{\mathbf{z}}\sim p(\mathbf{z})$ as illustrated in Fig. 2. Here, the likelihood and prior distributions are respectively modelled as
@@
The parameterization function for the mean $\mu_{\psi}$ can be chosen without a constraint and often modelled using a neural network $f^{\mu}_{\psi}(\mathbf{z},\mathbf{c})$. However, the parameterization of the covariance requires satisfying positive definiteness, and common practice is to model it as a positive diagonal matrix. Unfortunately, this choice results in an independence between the components of generated samples and hinders the modelling power.
@@ IV-A Dataset
This study uses a dataset of smart meter data collected from electricity consumers across 47 provinces in Spain, including homes, offices, and businesses [9]. The dataset spans from November 2014 to June 2022 and includes hourly electricity consumption records for 25,559 users. For the experiments, a subset was extracted focusing on data from Gipuzkoa, the province with the highest data density, between June 2021 and June 2022. Only users with at least one full year of continuous enrollment during this period were included, while those with negative and consistently zero consumption were excluded. The final dataset comprises $U$=6,830 users, each with $N$=365 daily profiles (24 hourly values per day, $T$=24), amounting to approximately 2.5 million records.
Since the data is zero-inflated, we used the same hyperparameters to apply a zero-preserved log-normalization to each feature as described in [6]. After normalization, we reserved the last 73 days of all users’ records (20% of the whole dataset) as the testing set. At the same time, the remaining portion was split into training and validation sets with a ratio of 3:1. We used the validation set for early stopping and learning rate scheduling of every model we trained.
We conduct the experiments using two auxiliary conditioning features: months and weekdays. These conditions are derived from the data index $n$, transformed using a cyclic ($\sin$-$\cos$) encoding as proposed in [10], and associated with their respective data points $\mathbf{x}_{un}$. As for the user embeddings, we used the same hyperparameter setting employed in [6].
@@ IV-B Baseline Model
$\begin{split}y^{(t)}_{0}&=f_{\phi}^{(t,0)}(\mathbf{x},\mathbf{c}),~{}\forall t\\ y^{(t)}_{i}&=y^{(t)}_{0}+\sum_{j=1}^{i}f_{\phi}^{(t,j)}(\mathbf{x},\mathbf{c}),~{}\forall i\geq 1,\forall t\end{split}$ (6)
where $y^{(t)}_{i}$ represents the $q^{(i)}$-th quantile value of the $t$-th feature for a given regressor $\mathbf{x}$ and contextual information $\mathbf{c}$ mapped using a neural network $f_{\phi}:\mathbb{R}^{T+C}\rightarrow\mathbb{R}^{T\times Q}$ with the constraint $f_{\phi}^{(t,i)}>0,\forall i\geq 1,\forall t$. This model can be trained by minimizing the following loss function
@@
for a given dataset $\{(\mathbf{y}_{m},\mathbf{x}_{m},\mathbf{c}_{m})\}_{m=1}^{M}$ and predictions $\{[\mathbf{y}_{mi}]_{i=1}^{Q}=[[y^{(t)}_{mi}]_{t=1}^{T}]_{i=1}^{Q}\}_{m=1}^{M}$. Note that this corresponds to the conventional quantile loss averaged over all output features and quantile levels. This setup can easily be converted to the day-ahead forecasting task at hand by replacing $\mathbf{y}_{m}$ with $\mathbf{x}_{un}$, $\mathbf{x}_{m}$ with $\mathbf{x}_{u(n-1)}$, and $\mathbf{c}_{m}$ with $\mathbf{c}_{un}$.
@@ IV-C Training
We applied the same training, neural network, and constraint settings on GUIDE-VAE as in [6], including features like the number of layers and neurons, learning rate, early stopping, and learning rate schedule. Also, we set the size of the quantile regression baseline model equal to the GUIDE-VAE in terms of number of learnable parameters to keep the comparison fair.
@@ IV-D1 Benchmarking Metrics
These metrics are used to test the performance of GUIDE-VAE against the quantile regression model. Since the quantile regression model does not yield an explicit probability distribution against which to compare, we adapted GUIDE-VAE to the quantile forecasting setting. For this, we took $S$ samples from $p_{\psi,\mathbf{U}}(\mathbf{x}_{un}|\mathbf{c}_{un})$ for each ($u$,$n$)-pair as $\hat{\mathbf{x}}_{uns}\sim p_{\psi,\mathbf{U}}(\mathbf{x}_{un}|\hat{\mathbf{z}}_{s},\mathbf{c}_{un}),\forall s\in[S]$ where $\hat{\mathbf{z}}_{s}\sim p(\mathbf{z})$. Recall that each condition $\mathbf{c}_{un}=\left[\mathbf{x}_{u(n-1)},\hat{\theta}_{u},\tilde{\mathbf{c}}_{n}\right]$ represents the look-back window and the contextual information required for forecasting $\mathbf{x}_{un}$. Then, we calculated the quantile values of each feature $x_{un}^{(t)}$ for given quantile levels $\mathbf{q}=[q^{(i)}]$ using these $S$ samples as
@@
Having access to quantile values for both forecasting models, we enlist the benchmarking metrics and their definitions:
@@ IV-D2 Ablation Metrics
    $\text{ALL}=\frac{1}{UN_{\text{test}}}\sum_{u,n}^{U,N_{\text{test}}}\log p_{\psi,\mathbf{U}}(\mathbf{x}_{un}|\mathbf{c}_{un})$ (11)
@@ VI Conclusion
This study introduced a probabilistic day-ahead forecasting framework based on GUIDE-VAE, a conditional VAE model designed for multi-entity datasets. GUIDE-VAE incorporates key features such as a pattern dictionary and entity embeddings, enabling efficient modelling of complex data structures. Unlike traditional approaches, GUIDE-VAE supports scalable, entity-specific probabilistic forecasting with a single model, eliminating the inefficiency of training separate models for individual entities. Its flexible outputs range from interpretable point estimates to full probability distributions that capture uncertainty and temporal dependencies.
Household electricity consumption was used as a challenging case study due to its highly stochastic nature and large number of entities. Our results demonstrate that GUIDE-VAE significantly outperforms conventional quantile regression techniques across key metrics. Additionally, the ablation study highlights the critical role of PDCC in capturing temporal correlations and improving forecast realism and accuracy across diverse settings.
The training phase is the most computationally demanding aspect of the GUIDE-VAE forecaster. It consists of two stages: entity embedding extraction and VAE training, where the latter is most demanding, and employing pattern dictionaries does not dramatically affect the VAE training time. In deployment scenarios, retraining the model, e.g., once a month, is not too burdensome. We note that the independent nature of the sampled profiles makes GUIDE-VAE’s inference easily parallelizable, making it fast in practice.
In conclusion, GUIDE-VAE offers a powerful and generalizable framework for probabilistic forecasting tasks in energy systems and beyond. Its ability to handle multi-entity data, provide entity-specific forecasts, and deliver probabilistic outputs makes it a versatile tool for decision-makers aiming to improve risk management, resource allocation, and operational planning. Future work will focus on extending GUIDE-VAE to other multi-entity datasets, such as feeders and wind turbines, exploring the use of its covariance structures for refining intra-day forecasts and uncovering deeper insights into temporal dependencies.